### v1

# PAID: A Framework of Product-Centric Advertising Image Design

Abstract.
In E-commerce platforms, a full advertising image is composed of a background image and marketing taglines. Automatic ad image design reduces human costs and plays a crucial role. For the convenience of users, a novel automatic framework named Product-Centric Advertising Image Design (PAID) is proposed in this work. PAID takes the product foreground image, required taglines, and target size as input and creates an ad image automatically. PAID consists of four sequential stages: prompt generation, layout generation, background image generation, and graphics rendering. Different expert models are trained to conduct these sub-tasks. A visual language model (VLM) based prompt generation model is leveraged to produce a product-matching background prompt. The layout generation model jointly predicts text and image layout according to the background prompt, product, and taglines to achieve the best harmony. An SDXL-based layout-controlled inpainting model is trained to generate an aesthetic background image. Previous ad image design methods take a background image as input and then predict the layout of taglines, which limits the spatial layout due to fixed image content. Innovatively, our PAID adjusts the stages to produce an unrestricted layout. To complete the PAID framework, we created two high-quality datasets, PITA and PIL. Extensive experimental results show that PAID creates more visually pleasing advertising images than previous methods.
Figure 1. Generated advertising images by our methods with only a product subject image, a target size, and taglines as input.
1. Introduction
Advertising images for products play a crucial role in commercial recommendation systems. Visually harmonious, aesthetically pleasing, and attractive advertising images tend to achieve relatively high click-through rates (Wang et al., 2021). With the rapid advancement of generative methods (Goodfellow et al., 2014; Menick and Kalchbrenner, 2019; van den Oord et al., 2016; Ho et al., 2020; Rombach et al., 2022a), it has become possible to automatically generate advertising images by providing a product foreground image, marketing taglines, and target size as input.
To accomplish this task, by referring to existing approaches (Du et al., 2024; Weng et al., 2024; Wang et al., 2024b; Lin et al., 2023; Zhou et al., 2022), we can first utilize heuristics and predefined rules to generate background text prompts. These prompts are then fed into text-to-image inpainting models to generate backgrounds for product foregrounds, resulting in a composite advertisement image. If marketing taglines are required, the layout of these taglines is predicted based on the generated image, and the taglines are rendered onto the image using techniques like attribute prediction. However, this approach suffers from several limitations: a) it cannot adaptively generate background prompts based on the specific content and shape of the product foreground, leading to poor foreground-background harmony and limited diversity; b) the position of the product foreground is fixed and cannot be adjusted according to the descriptions of the product foreground and background prompt, potentially resulting in inappropriate product size and poor composition; and c) when taglines need to be added, the image content is already finalized, restricting the spatial layout available for tagline placement.
Figure 2. Pipeline of PAID. It consists of four stages and generates advertising images centered around product information.
In this paper, we introduce a novel framework called Product-Centric Advertising Image Design (PAID). This framework centers around the integration of product foreground images and marketing taglines to automatically create more harmonious and aesthetic advertising images. Although PAID remains a multistage approach for generating advertising images, we have adjusted the stages and task definitions to optimize existing methods concerning product-background compatibility and the spatial arrangement of product foregrounds and marketing taglines.
As illustrated in Figure 2, we first introduce a task to generate prompts, using a Visual Language Model (VLM) (Dong et al., 2024) to adaptively create foreground and background prompts that better match the scene, placement, and angle based on product foregrounds. Following this, we propose a layout generation task that organizes the arrangement of the graphic elements (marketing taglines and decorative underlays), product foreground, and other nongraphic elements according to the product foreground, background prompt, marketing taglines, and target size. By advancing the layout stage and arranging the product foreground and marketing information concurrently, as opposed to generating the overall image first and then deciding the marketing text layout, the latter becomes less constrained by the image. This allows for a more rational presentation that is suitable for the needs of the content. In this process, the size and position of the product and other nongraphic elements are considered together, which addresses the limitations of adapting the foreground based on prompts in the inpainting task and reduces conflicts between nongraphic elements and marketing information that could adversely affect readability and attractiveness. Next, we train an inpainting text-to-image model based on SDXL (Podell et al., 2024) that can be controlled by input layouts. Combining inpainting, layout, and prompt control effectively is challenging due to potential conflicts among multiple control signals. A LoRA (Hu et al., 2022) adaption strategy is adopted to balance conflicts and facilitate training. Since layout mainly influences higher-level abstract information, we apply layout control only in the deeper blocks of the UNet, combined with a multi-scale training strategy to enhance training stability and efficiency. Finally, we utilize an attribute prediction and graphics rendering module to overlay graphic elements onto the image. To build the framework and verify its effectiveness, we have established a corresponding dataset, collecting a total of 38,017 images with detailed image captions, product foreground images, and layouts of graphic and nongraphic elements.
In summary, our main contributions are as follows:
  • We propose a novel multi-stage framework for advertising image design, which centers on product foregrounds and taglines and enhances the overall content prominence while maintaining the harmony and aesthetics of the images.
  • A VLM is introduced to generate suitable prompts for image generation using the input product foreground image, considering product placement angle and shape.
  • To arrange product foregrounds and taglines more effectively, a VLM-based layout generator is proposed. The generator can mix and arrange graphic and nongraphic elements with the target size and prompt. Accordingly, an image generation model that simultaneously accepts layout, prompt, and foreground control is proposed. Several strategies are introduced to enhance effectiveness and efficiency.
  • We created two datasets: one with 38,017 annotated images for advertising image design and another with 12,247 annotated images for background generation. We use these datasets to train and validate the models in the framework.
2. Related Work
2.1. Advertising Image Design
In recent years, various algorithms are proposed to automate advertising image design in e-commerce scenario (Zhou et al., 2022; Hsu et al., 2023; Li et al., 2023b; Seol et al., 2024; Lin et al., 2023; Li et al., 2023a). Typically, these approaches can be divided into two subcategories: background image-based and foreground-based methods. Background image-based methods adopt a full image as input and design models to predict the layout, followed by attribute prediction and graphics rendering techniques to construct the final ad image. Here, a layout means the class and position of texts, underlays, or other graphic elements. CGL-GAN (Zhou et al., 2022) and DS-GAN (Hsu et al., 2023) design transformer architectures to leverage the content of the background image and its saliency map for generating layout. RADM (Li et al., 2023b) builds a diffusion model and uses the image content and tagline content to derive the layout. It is the first work to consider tagline content. PosterLlama (Seol et al., 2024) leverages the rich design knowledge in Large Language Models (LLMs) to predict the layout. AutoPoster (Lin et al., 2023) introduces a complete pipeline which first generates a layout based on a background image, and then predicts the attributes of graphic elements and renders them onto the image. For these methods, the spatial layout is restricted due to fixed image content, which further restricts the tagline content, such as number and length.
In contrast, P&R (Li et al., 2023a) is a foreground-based method that leverages both the product foreground image and tagline content to predict the layout of the product foreground and taglines. It subsequently designs a module network to outpaint the background and render the texts according to the given background prompt and generated layout. The tagline content is not limited, and the layout is more flexible. However, P&R ignores the harmony between the background prompt and the generated layout. The layout may conflict with background generation. In our PAID framework, the overall layout of graphic elements, product foreground, and other nongraphic elements is generated jointly according to the product foreground, background prompt, and tagline contents, aiming to achieve the best harmony among them.
2.2. Controllable Image Generation
Diffusion models (Ho et al., 2020; Rombach et al., 2022a) have significantly improved image generation quality. To control instances and improve image quality for text-to-image models, some methods introduce layout guidance. These layout-to-image methods can be grouped into training-free and training-based categories. Mostly, training-free layout-to-image methods either manipulate the attention map of instances in prompt (Kim et al., 2023), or compute attention-based loss and back-propagate the gradients in de-noising procedure (Chefer et al., 2023; Xie et al., 2023; Chen et al., 2024). MultiDiffusion (Bar-Tal et al., 2023) conducts inference for each instance and then combines their latent vectors. Training-based methods modify the network architectures and fine-tune the diffusion model with layout-image data (Li et al., 2023c; Wang et al., 2024a; Zhou et al., 2024). GLIGEN (Li et al., 2023c) and InstanceDiffusion (Wang et al., 2024a) add a gated self-attention layer to introduce layout information. MIGC (Zhou et al., 2024) conducts per-instance shading and aggregates all the shaded instances.
For advertising image generation (Du et al., 2024; Wang et al., 2023; Cao et al., 2024; Li et al., 2023a), text-to-image models equipped with inpainting capabilities are frequently utilized. The inpainting technique is employed to preserve product characteristics, while the text prompt is used to describe the background. The foreground image is utilized by incorporating extra channels into the UNet architecture (Rombach et al., 2022a) or employing an inpainting ControlNet (Zhang et al., 2023). To better control the overall layout of advertising images and improve the quality of generated images, we combine inpainting with layout control. Currently, there is limited exploration in this area. SceneBooth (Chai et al., 2025) is the first model to integrate both controls in training, combining gated self-attention layers and inpainting ControlNet to train a model based on SD 1.5 (Rombach et al., 2022b) that can accept layout and foreground controls. However, directly combining the two to train a larger model like SDXL can lead to non-convergence or a decline in the quality of generated images. Therefore, we use several strategies when training the layout-controlled inpainting model to ensure training stability, generation quality, and efficiency.
3. PITA Dataset and PIL Dataset
For automatic product-centric advertising image design, we collect a large-scale Product-Centric Image-Text Advertising (PITA) dataset, gathering 38,017 samples from e-commerce platforms and the CGL dataset (Zhou et al., 2022). 1,000 of them are used for testing. PITA is more diverse than CGL, featuring images with various aspect ratios: 0.684 (51%), 1.0 (33%), 0.667 (14%), and 0.75 (2%). Each image is labeled with prompts (foreground caption and background caption), a product mask, a layout of graphic and nongraphic elements. Each element is represented with a type and a bounding box (bbox). Graphic elements contain “Logo”, “Text”, and “Underlay”. Figure 3 shows some examples in the PITA dataset.
Figure 3. Examples from the PITA dataset, with text, underlay, and logo highlighted in orange, yellow, and red rectangles, respectively.
For layout-controlled background inpainting, we additionally construct a Product-Centric Image Layout (PIL) dataset. The images do not contain graphic elements and labeled with prompts, product masks, and layouts of nongraphic elements. Figure 4 gives some examples. The image scenes cover indoors and outdoors. Product types include furniture, food, clothes, shoes, and so on. PIL contains 12,247 samples in total and 1,000 of them are divided for test.
Figure 4. Examples in PIL dataset.
4. Method
4.1. Overall Framework
As illustrated in Figure 2, our proposed PAID framework consists of four stages: prompt generation, layout generation, background image generation, and graphics rendering.
4.2. Prompt generation
For prompt generation, we fine-tune a large vision-language model $\pi_{prompt}$ based on InternLM-XComposer2-vl (XCP2) (Dong et al., 2024). The method is illustrated in the upper part of Figure 5.
Input data format. For the input of LLM, we first describe the prompt generation task and ask for foreground and background descriptions. For the image input, we crop out the product foreground and feed it into the visual encoder.
Output data format. To enable downstream layout and background image generation tasks, and to better understand the relationship between the foreground and background, we use the prompt generation model $\pi$ to simultaneously predict the foreground description $p_{fore}$ and the background prediction $p_{back}$. We format the output in JSON as it is compatible with the pre-trained VLM and simplifies the following analysis. The output JSON is structured as $p=(p_{fore},p_{back})$.
Training scheme. A standard VLM training scheme involves two stages: pre-training and supervised fine-tuning (SFT) (Liu et al., 2023a). During pre-training, the projector learns to align features between visual and text modalities. In the SFT stage, the projector and LLM are further trained for visual understanding tasks. Since we utilize the VLM trained after the SFT stage, the model can understand input images and follow instructions. We only fine-tune the LLM branch with our dataset to adapt the model to the prompt generation task, using cross-entropy loss as the objective function. The model $\pi_{prompt}$ takes the foreground image $I_{fore}$ and predefined instructions as input, performing two main tasks: generating a foreground description $p_{fore}$ and predicting the background description $p_{back}$. The prompt generation process is as follows:
(1) $p=\pi_{prompt}(I_{fore})$
4.3. Layout generation
In this stage, we introduce the Jointly Predict Graphic and Nongraphic Layout (JPGNL) method, which aims to optimize the arrangement and improve the readability and appeal of advertising images. It predicts the layout of graphic elements, the product, and other nongraphic elements based on the foreground image, prompt, provided taglines, and target size. Similar to prompt generation, we use XCP2 (Dong et al., 2024) for layout generation. The method is shown in the lower part of Figure 5.
Figure 5. The framework of our prompt and layout generation model and data format.
Input data format. For the input to the LLM, we describe the layout generation task and organize input conditions in JSON format. The image input remains the same as in prompt generation. A target canvas aspect ratio is given as condition to support multi-scale layout design. The aspect ratio of the product foreground is provided to help keep its shape. Foreground and background prompts are also given for generating the image layout. For the pre-defined taglines, we list them with empty locations and expect the model to complete them. If a logo is needed, we include it as a condition in the graphic layout section and specify its aspect ratio.
Except for these basic conditions, we also take product characteristics into account. In E-commerce scenarios, some products should not be occluded for better display. While for others, like clothing or human-hold items, partial occlusion is acceptable. To differentiate between these, we introduce a Class-Conditioned Layout Prediction (CCLP) strategy. CCLP provides clear information about the product class and whether foreground occlusion is allowed. For the not-allowing occlusion set (No Occ Set), we describe “The class of subject is [V1]. The bounding boxes of texts should never occlude the subject.” For the allowing occlusion set (Allow Occ Set), the description is “The class of subject is [V2]. The bounding boxes of texts are allowed to occlude the subject.”
Output data format. The output layout is structured in JSON format with three parts: subject layout, other non-graphic layout, and graphic layout. To maintain the shape of the product foreground, we propose a Ratio-Keeping Bbox Representation (RKBR). It is expressed as [x, y, h*r1/r2, h]. (x,y) is the normalized center point. w and h are the normalized width and height. r1 and r2 are the actual aspect ratio of the foreground and canvas respectively. Each value ranges from 0 to 1 with three decimal places. The nongraphic layout contains the location of elements in the background prompt, each having a type (instance name) and bounding box (bbox). The graphic layout defines each element by type, content, and bounding box, covering text, underlay, and logo. The content key is used only for text types to specify its semantic content. Bounding boxes are listed as [x, y, w, h]. The model is expected to supplement underlay adaptively. If logos are required, it is represented in the same way as the product for ratio keeping.
Training scheme. Same as the prompt generation module, we only fine-tune the LLM branch with the visual encoder and projector frozen. The output JSON is structured to predict the layout in a raster scan order, adhering to the ascending order of top-left coordinates. Furthermore, we shuffle the taglines in the input conditional JSON for data augmentation and preventing the model from memorizing the input sequence.
4.4. Background Image generation
In this process, we train an SDXL-based (Podell et al., 2024) image generation model controlled by product foreground, prompt, and layout.
Layout control. Suppose there are $N$ elements to be drawn in an image, represented as $\{(n_{i},l_{i})|i=0,1,...,N-1\}$. We map these into $N$ embeddings as layout input. Here, $n_{i}$ is the element name, and $l_{i}$ is its location. Concretely, we use the text encoder in SDXL to map $n_{i}$ into a global semantic embedding $\textbf{e}_{n,i}$. $l_{i}$ is transferred into normalized top-left and bottom-right coordinates, then fed into a Fourier mapping layer(Tancik et al., 2020), deriving the location embedding $\textbf{e}_{l,i}$. Next, $\textbf{e}_{n,i}$ and $\textbf{e}_{l,i}$ are concatenated and fed into a trainable MLP layer to represent the layout information of the $i$-th element, denoted as $\textbf{e}_{i}$. $\textbf{e}=\{\textbf{e}_{i}|i=0,1,...,N-1\}$ is the layout embeddings. A gated self-attention (GSA) layer (Li et al., 2023c) is inserted between the self- and cross-attention layers in the original UNet to inject layout embeddings. Supposing there are $M$ visual embeddings (V) after the self-attention layer. The output of the GSA layer is
$GSA(\textbf{V},\textbf{E})=\textbf{V}+tanh(\gamma)*SA(cat(\textbf{V},\textbf{E}))[:M]$
, where $\gamma$ is a learnable parameter, $cat$ means concatenation, and $SA$ means self-attention.
Considering that layout is semantic-level information and less related to high-frequency details, we only add GSA layers in the deep blocks of UNet. We refer to this strategy as Deep Layer Control (DLC). Specifically, the layout control is applied in the middle block and the lowest resolution of up blocks. This reduces the parameters and inference latency, and has no negative impact on generation quality. Additionally, we adopt a LoRA Adaptation Training (LAT) strategy that trainable LoRA (Hu et al., 2022) layers are added to UNet, for bridging the gap between the base model and added GSA layers. This further accelerates convergence and increases generation quality. Figure 6 describes how layout control is achieved.
Figure 6. Illustration of layout control.
Training stages. To build the product-reserving layout-to-image model, the training involves two steps. First, we modify the UNet architecture in SDXL by adding the above GSA layers and train it for deriving a layout-to-image model. We use a multi-scale training strategy for faster convergency, first training on 512x512 resolution images and then fine-tuning on 1024x1024 resolution images. Second, we equip the trained model with an inpainting ControlNet and further fine-tune it on 1024x1024 resolution images. This step incorporates the product foreground, achieving full control over the prompt, layout, and product foreground.
4.5. Graphics Rendering
Since there is no effective method to predict the visual attributes of text and other graphic elements, and the graphics rendering module is not the focus of this paper, we design a strategy according to the aesthetic rules to handle graphics rendering. We first sort the text elements by area to decide the font and color, both of which are limited to a specific set (Liu et al., 2024). The selection is based on the similarity to the foreground color and contrast with the background area. Texts grouped by size and position use the same font and color. For the underlay, the color is chosen based on its contrast with the text color and similarity to the background color. The shape is selected from a predefined SVG library based on its size and proportions, with minor adjustments for fitting.
5. Experiments
In this section, we compare our PAID framework and its modules with previous ones. We conducted experiments on the PITA dataset, and the results confirm the effectiveness of our method.
5.1. Implementation Details
For prompt and layout generation, XCP2 (Dong et al., 2024) is used as our backbone. Its visual encoder is built upon CLIP ViT-Large (Radford et al., 2021), processing 490x490 resolution images. The LLM is based on InternLM2 (Cai et al., 2024). We fine-tune the LLM branch on the PITA dataset for one epoch. We train the model using AdamW optimizer with a learning rate of 1e-5. Batch size on each GPU is set to 1 and gradient accumulation of 8 is adopted. The training costs about 14 hours with 8 NVIDIA A100 GPUs. During inference, we use top-p sampling with p set to 0.9 and a sampling temperature of 0.6.
For background image generation, we use SDXL base model as the backbone. During the first step, about 6.8 million high-aesthetic internal images with layout annotations are used. We train the newly added layers with a learning rate of 5e-5 for 20k steps on 512x512 resolution and further fine-tune them with a learning rate of 2e-5 for 20k steps on 1024x1024 resolution. Then in step 2, we introduce an inpainting ControlNet which is pre-trained with millions of advertising images from E-commerce platforms (Creative, 2024), and fine-tune the layout-related layers with a learning rate of 2e-5 for 1k steps, using 12k training data in the collected PIL dataset.
5.2. Evaluation Metrics
Overall pipeline. We evaluate the effectiveness of the overall pipeline from two perspectives: overall visual quality and layout. For visual quality, we employ the Fréchet Inception Distance (FID) (Heusel et al., 2017) and aesthetic score. The aesthetic score is obtained by advertising experts ranking the outputs of different pipelines. The average rank is calculated to represent the mean aesthetic rating, where a lower rank indicates superior creative results. For layout, we adopt the metrics used in previous studies (Zhou et al., 2022; Seol et al., 2024), which are detailed in the subsequent Layout Generation module.
Prompt generation. We evaluated prompt generation quality based on fore-background matching and e-commerce domain score.
  • FBM Rate (Fore-background matching rate): It refers to how well a product foreground fits with a given prompt background. We classify this matching into “reasonable” or “unreasonable” through human annotation, and then determine the proportion of reasonable cases for each method.
  • ED Score (E-commerce domain score): To assess how well the model-generated prompts reflect e-commerce characteristics, we calculate the FID between the CLIP features of these prompts and the ground truth prompts in the test set.
Layout generation. Following previous work (Zhou et al., 2022; Seol et al., 2024), we use graphic metrics and content metrics to evaluate layouts. Graphic metrics focus on relationships between graphic elements, including validity $Val$, alignment $Ali$, overlap $Ove$, and underlay $Und_{l},Und_{s}$. Val means the ratio of elements larger than 0.1% of the canvas. Ali assesses how well elements are spatially aligned. Ove is the average intersection over union (IoU) of all element pairs except for underlays. $Und_{l}$ and $Und_{s}$ checks if underlays enhance non-underlay elements. Content metrics assess the harmony of layout with the background image, including utility $Uti$, occlusion $Occ$, and unreadability $Rea$. Uti evaluates space usage for graphic elements. Occ is the average overlapping area between the graphic elements and products. Rea represents the non-flatness of regions that text elements without underlays are put on. For methods which take taglines as input, we report the text match rate (TMR), to see if the number of generated text bboxes are consistent with input taglines.
Background image generation. We evaluate the quality of generated background images with FID, CLIP-T (Ruiz et al., 2023), CLIP-I (Gal et al., 2023). To verify layout control, we use Grounding-DINO (Liu et al., 2023b) to detect instances and calculate the maximum IoU between detected bboxes and ground-truth. The average of these maximum IoUs is called mIOU. In addition, Grounding-DINO AP scores are also reported.
5.3. Comparison
Compare with automatic pipelines for product advertising image design. Since there are no public automatic methods for advertising image design conditioned on the product foreground, taglines and target size, we design two pipelines for comparison. Pipeline 1 use GPT-4o (OpenAI, 2024) for prompt and layout generation based on the input conditions. We fine-tune the SDXL-based inpainting model on our PIL dataset to obtain the image generation model. The only difference between this method and ours is the absence of layout control and the corresponding training stage. During inference, only the product foreground is controlled by the layout output of GPT-4o. Pipeline 2 first generate a background prompt according to the foreground image using GPT-4o. The image generation model is the same as in Pipeline 1, but during inference, the foreground is placed according to general athetic rules (see details in the Appendix). Next, the image and taglines are processed by the SOTA layout generation model PosterLlama to generate a graphic layout. Last, the results of these two pipelines are processed through the same graphics rendering module as our method to produce the final image. Table 1 shows the quantitative comparison between these pipelines and our PAID framework. PAID outperforms in most metrics, except for slightly higher occlusion compared to Pipeline 1, which places the product centrally with a small bounding box. However, Pipeline 1 uses GPT-4o for layout, lacking product-centric design knowledge, resulting in lower quality. PAID uses trained experts for different tasks in ad design, creating more visually appealing results (see Figure 7). Pipeline 2 fixes the product location, limiting adaptability to product shapes, whereas PAID adjusts layouts based on product shape for better visuals. Additionally, the fixed background of Pipeline 2 can lead to layouts with crowded text boxes, making them hard to read, while our model provides more flexible and readable layouts.
Table 1. Quantitative comparison between automatic pipelines for product-centric advertising image design.
Overall Visual Quality Layout Quality
[…]
Pipeline 1 56.545 2.781 0.0021 0.002 0.8333 0.7872 0.0891 0.0937 0.234
Pipeline 2 43.221 1.814 0.0016 0.0028 0.9994 0.9930 0.0984 0.1158 0.1864
PAID 37.524 1.405 0.0013 0.0017 0.9999 0.9973 0.1367 0.0955 0.1815
Figure 7. Visualization of advertising image designed by different methods.
Compare with prompt generation methods. For prompt comparison, we use GPT-4o and an untrained version of our model (XCP2) as baselines. Both utilize in-context learning (ICL) for optimal performance. The FBM Rate is determined by votes from three annotators for each case. As shown in Table 2, our method outperforms the baselines in both FBM Rate and ED Score. We also use t-SNE for dimensionality reduction to visualize the feature distribution of prompts generated by our model before and after training (see Figure 8). This shows that our model effectively aligns with the e-commerce prompt dataset distribution after training.
Table 2. Evaluation of prompt generation methods
Method FBM Rate $\uparrow$ ED Score $\downarrow$
GPT-4o 97.9% 22.49
XCP2 94.8% 13.04
Finetuned XCP2 98.5% 7.17
Figure 8. T-sne visualization of generated prompts.
Compare with layout generation methods. To verify the advantage of our layout generation model, we compare it with previous background-based methods (DS-GAN (Hsu et al., 2023), RADM (Li et al., 2023b), and PosterLlama (Seol et al., 2024)) and the foreground-based method P&R (Li et al., 2023a). For background-based generation methods, we conduct experiments on two versions of the testing set for fair comparison, named the erased set and the generated set. The erased set contains images with graphic elements erased by LaMa (Suvorov et al., 2022), which may provide hints to put graphic elements on these erased area. The generated set contain images re-generated using their ground-truth prompts (image captions). The generated set has no above hints and conforms to real applications. As for the foreground-based method P&R, since it has not released the codes, we re-implement its core idea: first, generating the layout of the foreground and graphic elements conditioned on the foreground and taglines, and second, generating the final image based on the background prompt, layout, and product foreground. Our method uses the ground-truth prompt, product foreground, and taglines in PITA for prediction. The results are shown in Table 3 and Figure 9. Background-based methods perform relatively poorly in terms of product occlusion and overlap and alignment of graphic elements. This may be due to the fixed background image content, leaving less room for the model flexibility. RADM and DS-GAN, which are insensitive to tagline content, also face issues with generating text boxes that cannot adequately fill the tagline content. The foreground-based method P&R improves in these areas but does not consider the harmony between the product, background prompt, and graphic layout, sometimes leading to unreasonable generated images. Our method, PAID, takes into account the shape of the product foreground, background prompt, and tagline content to decide the overall layout, including graphic and nongraphic elements, resulting in a better overall effect.
Table 3. Quantitative performance comparison of layout generation methods on PITA dataset. Bold number represents the best result in each column except for erased set. BG-based and FG-based are the abbreviations for Background-based and Foreground-based methods, respectively.
[…]
BG-based (Erased) DS-GAN(Hsu et al., 2023) 0.9585 0.0270 0.0058 0.3910 0.0744 0.1816 0.1063 0.1826 -
RADM(Li et al., 2023b) 0.999 0.0411 0.0017 0.9852 0.6934 0.1484 0.0814 0.1722 0.82
PosterLlama(Seol et al., 2024) 0.9984 0.002 0.0026 0.9899 0.9870 0.1226 0.0976 0.1820 0.998
BG-based (Generated) DS-GAN(Hsu et al., 2023) 0.9621 0.0284 0.0080 0.3350 0.0528 0.1701 0.1202 0.2328 -
RADM(Li et al., 2023b) 0.9753 0.0484 0.0201 0.7997 0.2828 0.0657 0.2527 0.2558 0.289
PosterLlama(Seol et al., 2024) 1.0 0.0015 0.0022 0.9990 0.9965 0.1090 0.1208 0.231 0.996
FG-based P&R(Li et al., 2023a) 1.0 0.0012 0.0019 0.9966 0.9929 0.1367 0.1000 0.2008 1.0
Ours 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968 1.0
Figure 9. Qualitative comparison of layout generation methods on PITA test set. Orange: text, Yellow: underlay, Red: logo, Black: Invalid
Compare with layout-controlled inpainting models. We compare our layout-controlled background image generation model with others, including training-based method GLIGEN (Li et al., 2023c) and training-free methods BoxDiff (Xie et al., 2023), TFCLG (Chen et al., 2024), and MultiDiffusion (Bar-Tal et al., 2023). Since these methods lack inpainting ability, we re-implement them with the SDXL-based inpainting ControlNet (Creative, 2024). Quantitative results are reported in Table 4. Training-free methods have weaker spatial control compared to training-based methods. Our trained model outperforms others in FID, mIoU, and AP, indicating better image quality and spatial control. As shown in Figure 10, other models tend to miss some instances when more objects are required, while our model generates them more accurately.
Figure 10. Visualization of images generated by different layout-controlled inpainting models.
Table 4. Quantitative performance comparison of layout-controlled inpainting models on PIL test set.
[…]
BoxDiff(Xie et al., 2023) 29.383 0.315 0.896 0.555 0.037/0.064/0.033
TFCLG(Chen et al., 2024) 29.498 0.315 0.895 0.543 0.033/0.056/0.029
MultiDiffusion(Bar-Tal et al., 2023) 42.359 0.311 0.841 0.5617 0.029/0.052/0.025
GLIGEN(Li et al., 2023c) 27.237 0.312 0.898 0.561 0.047/0.073/0.044
Ours 25.917 0.313 0.906 0.705 0.079/0.127/0.078
5.4. Ablation Studies
Effect of JPGNL in layout generation. Our method predicts the overall layout of graphic and nongraphic elements. Here, we examine the need to predict the layout of nongraphic elements according to the background prompt. As shown in Figure 11, without specifying the position of nongraphic elements, texts placed on complex areas can lead to a messy appearance. By predicting the overall layout, we improve the text readability of advertising images. Additionally, with nongraphic layout predictions, generated images align more closely with background prompts. Without JPGNL, images might miss some objects from the prompts. Besides, in Table 5, we quantitatively analyze the impact of predicting nongraphic layouts on graphic layouts. Predicting nongraphic layouts (w/ JPGNL) can improve metrics related to product occlusion and text readability. However, since it adds complexity to the task, it negatively affects graphic metrics such as element overlap and alignment.
Figure 11. An illustration of the effect of JPGNL. Predicting image layout makes better visual effect.
Table 5. Quantitative ablation study on JPGNL.
[…]
w/o JPGNL 1.0 0.0009 0.0013 0.9995 0.9891 0.1377 0.1001 0.2003
w/ JPGNL 1.0 0.0012 0.0017 0.9976 0.9956 0.1364 0.0973 0.1968
Effect of RKBR and CCLP in layout generation. We validate the effectiveness of the proposed input and output format for layout generation, including RKBR and CCLP. As shown in Table 6, without RKBR, the model may predict a product size which does not match the original aspect ratio about 5% of the time. Note that we regard a difference of less than 1.5% as correctness. Adding CCLP improves the Uti and Occ values on No Occ Set, indicating that CCLP helps the model to distinguish occlusion-allowing sets and others by providing explicit class conditions.
Table 6. Ablation study on input/output format construction of layout generation. O, NO, and AO correspond to the overall set, the No Occ Set, and the Allow Occ Set, respectively. FRC represents Fg Ratio Correctness.
[…]
1.0 0.0012 0.0018 0.9982 0.9971 0.1353/0.1448/0.0876 0.0985/0.0230/0.4258 0.2005 0.955
1.0 0.0012 0.0018 0.9991 0.9912 0.1360/0.1466/0.0816 0.1054/0.0209/0.4491 0.1994 1.0
1.0 0.0008 0.0018 0.9989 0.9906 0.1366/0.1492/0.0851 0.1002/0.0156/0.4282 0.1973 0.957
1.0 0.0012 0.0017 0.9976 0.9956 0.1364/0.1493/0.0849 0.0973/0.0136/0.4316 0.1968 1.0
Effect of DLC and LAT in layout-to-image model. We investigate the effect of deep layer control (DLC) and Lora adaptation training (LAT), with results shown in Table 7. Adding layout control in deep UNet layers performs almost as well as controlling all layers. It reduces parameters and inference costs, which benefits the application. With LoRA adaptation, the model achieves lower FID and higher mIoU, enhancing image quality and spatial control. Parameters and inference costs can be further reduced.
Table 7. Quantitative performance comparison of layout-to-image methods on PIL dataset.
[…]
27.237 0.312 0.898 0.561 0.047/0.073/0.044
27.308 0.312 0.898 0.558 0.045/0.075/0.043
25.675 0.313 0.906 0.696 0.076/0.123/0.075
25.917 0.313 0.906 0.705 0.079/0.127/0.078
6. Conclusion
In this paper, a novel framework named Product-Centric Advertising Image Design (PAID) is proposed to automatically create ad images only based on product foreground images, marketing taglines, and target sizes. PAID consists of four serial stages: prompt generation, layout generation, background image generation, and graphics rendering. First, it generates adaptive prompts according to the semantic and visual information of the product, for the benefit of background image generation. Then, it predicts how graphic and nongraphic elements should be placed according to the prompt, product foreground, and taglines, creating a harmonious overall layout. Next, a layout-controlled inpainting model is utilized for background image generation. A graphics rendering module is then applied to get the final images. We train separate experts to conduct these sub-tasks. Two high-aesthetic and diverse product-centric datasets are created for the convenience of training. Qualitative and quantitative comparisons prove that the proposed PAID framework generates more visually pleasing ad images, compared with previous methods.
Acknowledgements
This work is supported by National Natural Science Foundation of China (NSFC No.62076031 and No.62076036) and Alibaba Research Intern Program.